For long term deposits, Blueberry Bank pays 6.5% per annum, while Yellowbrook Bank pays 6.4%per annum compounded monthly. A bad decision will lead to your savings deposit of $10,000 for 5 years, being less by ________ (nearest dollar) in ??? bank.

On the other hand, for long term deposits, Savings Bank pays 6.5% per annum, while Bank of the North pays 6.25%per annum compounded monthly. A good decision will lead to your savings deposit of $10,000 for 5 years, having an extra __________ (nearest dollar) in ??? bank.

SHOW ALL WORK.
Answer #1

1.

Blueberry bank=10000*(1+6.5%)^5=13700.86663

Yellowbrook bank=10000*(1+6.4%/12)^(5*12)=13759.57275

A bad decision will lead to your savings deposit of $10,000 for 5 years, being less by 59 in Blueberry bank.

2.

Savings bank=10000*(1+6.5%)^5=13700.86663

Yellowbrook bank=10000*(1+6.25%/12)^(5*12)=13657.29902

A good decision will lead to your savings deposit of $10,000 for 5 years, having an extra 44 in Savings bank.